OpenStudy (anonymous):

how about this one 3x - 2y =11

OpenStudy (anonymous):

i got\[y = -\frac{ 1 }{ 2 }x - 5.5\] @tHe_FiZiCx99

OpenStudy (anonymous):

but its not right

OpenStudy (anonymous):

Are you looking for y?

OpenStudy (anonymous):

trying to get into slope intercept form

OpenStudy (anonymous):

y = 1/2 (3 x-11) That's for y.

OpenStudy (the_fizicx99):

3x - 2y = 11 -3x -3x -2y = -3x + 11 /-2 /-2 /-2 y = 3/2x - 5.5

OpenStudy (the_fizicx99):

Not 1/2x

OpenStudy (anonymous):

oh ok